OpenStudy (anonymous):

A system of equations is shown below. 5x - 5y = 10 3x - 2y = 2 Part A: Create an equivalent system of equations by replacing one equation with the sum of that equation and a multiple of the other. Show the steps to do this.

OpenStudy (anonymous):

Please help.

OpenStudy (yanasidlinskiy):

The Solution to the original problem is: x = -2 y = -4 part A: 5x-5y=10 (3x-2y=2) * 2 5x-5y=10 6x-4y=4 add: 11x-9y=14 New System: 11x-9y=14 3x-2y=2 Hope this helps!:)

OpenStudy (anonymous):

Yes but why did you multiply the second equation by two and not the other? Do you not have to do the same to both equations?

OpenStudy (anonymous):

I just want to figure out how to do this myself to I'm not looking for anyone to give me answers.

OpenStudy (yanasidlinskiy):

Do you know how to solve the original equation?

OpenStudy (anonymous):

Which one? There are two in the original system.

OpenStudy (yanasidlinskiy):

Either one.

OpenStudy (anonymous):

I can find the solution of them both, by getting them into y=mx+b form and then graphing them right?